# Training on the Test Task Confounds Evaluation and Emergence

### v1

@@ Abstract
We study a fundamental problem in the evaluation of large language models that we call training on the test task. Unlike wrongful practices like training on the test data, leakage, or data contamination, training on the test task is not a malpractice. Rather, the term describes a growing set of techniques to include task-relevant data in the pretraining stage of a language model. We demonstrate that training on the test task confounds both relative model evaluations and claims about emergent capabilities. We argue that the seeming superiority of one model family over another may be explained by a different degree of training on the test task. To this end, we propose an effective method to adjust for training on the test task by fine-tuning each model under comparison on the same task-relevant data before evaluation. We then show that instances of emergent behavior largely vanish once we adjust for training on the test task. This also applies to reported instances of emergent behavior that cannot be explained by the choice of evaluation metric. Our work promotes a new perspective on the evaluation of large language models with broad implications for benchmarking and the study of emergent capabilities.222Code and fine-tuned models are available at https://github.com/socialfoundations/training-on-the-test-task
@@ 1 Introduction
In this work, we group strategies to utilize task knowledge at training time under the umbrella term of training on the test task. Examples of training on the test task include the use of instruction-tuning data or question answering templates during pre-training (Zeng et al., 2022; Bai et al., 2023). We work from the premise that training on the test task is acceptable—or at least, unavoidable.
In a nutshell, we show that training on the test task strongly confounds model comparisons across different scales and model families. Moreover, it significantly obscures the study of emergent capabilities of large language models. Perhaps counterintuitively, we propose to mitigate the effects of training on the test task by doing more of it. We show that we can effectively level the playing field by giving each model the same, sufficient task-specific fine-tuning before evaluation. This adjustment restores clean log-linear scaling and makes capabilities predictable based on much smaller model scales.
@@ 1.1 Our contributions
We introduce the term training on the test task to group a growing repertoire of practices that utilize knowledge about evaluation tasks at training time. We study its impact on benchmark evaluations by inspecting 53 different language models in two major active benchmarks, MMLU and GSM8K.
@@
We start in Section 2 by dividing models into those trained before November 2023 and those trained after. We find that for the same amount of compute, newer models outperform older models on average by 7 percentage points in MMLU and 17 points in GSM8K. We then fine-tune all models on the same amount of task-specific data before evaluation. We show that after fine-tuning, newer models no longer outperform older models. See Figure 1. This outcome suggests that newer models outperform older ones because they—implicitly or explicitly—trained more on the test task. Moreover, it shows how test task training can distort benchmark performance.
@@
We propose a simple and effective method to adjust for the effect of training on the test task. Put simply, we fine-tune each model on the same, sufficient amount of task-specific data before evaluation. To validate our method, we demonstrate its effectiveness in a controlled setting: we take the older models and fine-tune half of them on the test task. Remarkably, this recreates the kind of performance differences observed between newer and older models. We then show that we can undo the advantage of the fine-tuned models over the other models by further fine-tuning all models on the test task (Section 3.1, Figure 3).
@@
Next, we give evidence that training on the test task may be a more dominant factor in benchmark performance than data contamination. To argue this point, we consider ARC and HellaSwag, which use cloze prompts for evaluation. Here, at first there appears to be no sign of an unfair advantage in any specific model family. But after switching to MMLU-style multiple choice prompts, we see the same confounded results as for MMLU (Section 3.2, Figure 4). This suggest that newer models perform well in MMLU likely not because of memorization of specific testing data, but rather due to an improved ability to comprehend MMLU-style prompts. Either way, our proposed adjustment recovers fair model comparisons.
@@
We show that training on the test task significantly distorts model family comparisons. The design choices of certain model families –such as its pretraining data mixture– may appear superior to others before adjusting for test task training but not after adjustment (Section 4.1, Figure 6). We also demonstrate that test task training overestimates the progress in capabilities achieved by recent models. After adjustment, newer models only modestly improve the Pareto frontier of performance against compute (Section 4.2, Figure 7).
@@
Finally, we demonstrate that training on the test task has profound implications for the study of emergent capabilities. Specifically, we show that the phenomenon of emergence disappears gradually as the amount of training on the test task grows (Section 5). In particular, we can make capabilities visible and predictable from much smaller model scales. Importantly, our adjustment also works in cases, like MMLU, where previous purported explanations of emergence, such as the choice of evaluation metric, do not suffice.
@@
Our work calls for a major reorientation of large language model evaluation. Model comparisons, scaling laws, and claims of emergence, are all strongly confounded by the choice of training data relative to the test tasks. Rather than scrambling to detect and disallow various forms of training on the test task, we propose to “fight fire with fire”. Specifically, our recommendation is to give each model the same sufficient amount of fine-tuning on task-relevant data prior to evaluation.
@@
Although deliberate training on the test task can level the playing field, our work shows that generally significant fine-tuning is needed before the playing field is actually level. This requirement poses additional computational burden on the side of the evaluator. The computational resources to fine-tune all models equally may not be available to all evaluators depending on the circumstances. In addition, sufficient task-relevant training data might be expensive to source or generally unavailable for many tasks. While this is not an issue when model trainers also lack access to such data, training on proprietary task-specific training data would be difficult to correct for.
@@ 2 Adjusting for training on the test task
We choose MMLU (Hendrycks et al., 2020) and GSM8K (Cobbe et al., 2021) as a case study for investigating training on the test task in active benchmarks. MMLU tests for world knowledge, whereas GSM8K tests multi-step mathematical reasoning. These two benchmarks are very prominent in the literature at present time. For instance, GPT 4 (Achiam et al., 2023), Claude 3 (Anthropic, 2024), Gemini (Gemini et al., 2023) and Llama 3 (MetaAI, 2024) all report and highlight MMLU and GSM8K. They are also included in the HuggingFace (HF) Open LLM Leaderboard111See Appendix C for results pertaining to the OpenLLM Leaderboard v2 (Fourrier et al., 2024a). (Beeching et al., 2023), a popular benchmark leaderboard that evaluates and ranks models with publicly available weights. We evaluate models using LM Evaluation Harness library (EleutherAI, 2024), in identical fashion to the HF leaderboard.
We evaluate 53 base models, ranging in size from 70M to 70B parameters. See Appendix A.1 for the full list. The HF leaderboard’s FAQ makes the distinction between “base pretrained models” and instruction-tuned or chat models, arguing that this is necessary to ensure fair model comparisons. We select models that are categorized as “pretrained”. We check that the technical report of each of the selected models makes no mention of the model being fine-tuned. We only consider models for which the number of training tokens is known. This allows us to estimate the total amount of pretraining compute in FLOPs as , where is pretraining compute, is the number of model parameters, and is the number of training tokens.
@@
We evaluate models on MMLU and GSM8K, and plot benchmark accuracy against pretraining compute in Figure 1 top. We observe that performance correlates with pretraining compute for both benchmarks. However, on the surface it appears that more recent models better leverage pretraining compute. In other words, for a given compute budget newer models are able to attain better benchmark performance. In fact, models trained after November 2023 Pareto dominate those trained before November 2023.
These improvements in benchmark performance coincide with a recent trend in LLM research of increasingly utilizing test task knowledge at training time. For example, Qwen 1.5 (Bai et al., 2023), Olmo 1.7 (Groeneveld et al., 2024) and MAP Neo (Zhang et al., 2024) include instruction data during pretraining. StableLM 2 (StabilityAI, 2023) reformulates some of its pretraining datasets to better resemble downstream tasks such as question-answering. More subtly, the pretraining data mixture of Gemma (Gemma et al., 2024) was determined partially based on downstream benchmark evaluations.
This raises an important question: Do newer models outperform older ones mainly because newer models effectively trained more on the test task? At first sight, an answer seems elusive. After all, it would be both infeasible and cost prohibitive to train all models with the same training data and compute budget. Nevertheless, in the next section, we propose a way to get at the answer by adjusting for the effect of training on the test task.
@@ 2.1 Adjusting for training on the test task by training on the test task
We propose to adjust for differences in test task training by fine-tuning all models on the same, sufficient amount of task-specific data before evaluation. To do so, we need a source of task-specific data for each of the tasks we consider. For multiple choice questioning answering (MMLU), we use the auxiliary training set accompanying the HF MMLU repository222https://huggingface.co/datasets/cais/mmlu. It contains around 100,000 training examples and around 30M tokens. For mathematical reasoning (GSM8K), we combine the the MetaMathQA (Yu et al., 2023b) and Orca-Math (Mitra et al., 2024) datasets, totalling 600,000 training examples and approximately 200M tokens. We fine-tune models for three epochs using standard hyperparameter choices, with minimal hyperparameter tuning, see Appendix A.2. Note that the amount of compute required for fine-tuning is minimal in comparison to the compute required for pretraining, since all models considered were pretrained on at least 300B tokens.
We plot model scores on MMLU and GSM8K after fine-tuning in Figure 1 (bottom). We observe that after fine-tuning on task relevant data, newer models no longer Pareto dominate in terms of accuracy per pretraining compute. Instead, benchmark performance is strongly correlated with compute and both newer and older models follow remarkably similar scaling trends. That is, newer models no longer appear to outperform older models. Moreover, we observe that older models tend to benefit from training on the test task much more than newer models, see Figure 2. The improvements of older models are striking, often jumping from random chance accuracy to double digit improvements in accuracy. In contrast, fine-tuning brings comparatively little benefit to newer models. This observation suggests that newer models have already been trained on a substantial amount of task-relevant data.
@@ 2.2 Quantifying performance differences between newer and older models
We draw inspiration from scaling laws (Kaplan et al., 2020) in how we model benchmark accuracy to scale log-linearly with pretraining compute . To account for emergence (Wei et al., 2022), we assume that models perform at the task’s random chance accuracy up to scaling to some point of emergence . We let the variable denote whether a model was trained after November 2023, and regress the model
where , and are the fit’s parameters, and is random noise. We focus on the coefficient , which corresponds to the average difference in benchmark performance between newer and older models after controlling for pretraining compute. We fit the model in Equation 1, and report the regression coefficient in Figure 1. We obtain R for all model fits. Before adjusting for test task training, the estimated difference in performance between newer and older models are statistically significant, positive, and large. Specifically, recent models on average outperform older ones by over 7 accuracy points in MMLU and 17 accuracy points in GSM8K. These are remarkable differences in benchmark performance, as small single digit improvements are typically considered substantial improvements by the literature.
We repeat the analysis but using models’ adjusted benchmark scores, that is, those obtained after fine-tuning on the test task. After adjusting for test task training we find no evidence for a significant difference in benchmark performance between newer and older models. That is, the estimated coefficient is both small and not statistically significant. Put simply, newer models no longer outperform older ones. Therefore, conditioned on all models training on the same, sufficient amount of task-specific data before evaluation, there are no differences in benchmark performance between newer and older models.
Our findings provide evidence that the differences in benchmark performance between newer and older models are largely attributable to differences in test task training. We present a causal interpretation of our results in Appendix B, outlying the causal assumptions needed to establish that the improvements of newer models are attributable to training on the test task. Overall, we find no evidence for the improvements in performance of newer models being attributable to anything other than training more on the test task.
We include in Appendix B.1 a robustness check on the temporal split chosen, by instead diving models based on whether they were trained primarily on English language data. We obtain similar differences in performance, which we interpret as a valuable robustness check of our results. In Appendix C we instead consider the benchmarks of the newly released HF OpenLLM Leaderboard v2 (Fourrier et al., 2024a). Whereas the HF leaderboard v2 pays particular attention to guarding against data contamination (Fourrier et al., 2024b), we nonetheless find evidence that training on the test task confounds all benchmarks included in the Leaderboard v2. These findings highlight that training on the test task is a distinct phenomenon from data contamination, and new methods –such as our proposed adjustment procedure– are required to mitigate the confounding effect of training on the test task on benchmark evaluations.
@@ 3 Recreating differences in benchmark performance
Previously, we introduced a way to adjust for training on the test task. Here we systematically test the validity of this adjustment method. To do so, we demonstrate how to recreate the observed differences in performance between newer and older models by actively manipulating how much models train on the test task.
We do so in two ways. First, we fine-tune older models on task relevant data (Section 3.1). Second, we reformulate certain test tasks to use multiple choice prompts instead “cloze” evaluation (Section 3.2). Both experiments turn out to recreate the kind of performance difference we observed earlier. This not only provides further evidence that differences in performance between older and newer models are linked to test task training. It also demonstrates how test task training distorts benchmark evaluations.
Fortunately, in both cases, we show that fine-tuning models on task-relevant data before evaluation is an effective mechanism for mitigating the bias introduced by training on the test task.

### v2

@@ Abstract
We study a fundamental problem in the evaluation of large language models that we call training on the test task. Unlike wrongful practices like training on the test data, leakage, or data contamination, training on the test task is not a malpractice. Rather, the term describes a growing set of practices that utilize knowledge about evaluation tasks at training time. We demonstrate that training on the test task confounds both relative model evaluations and claims about emergent capabilities. We argue that the seeming superiority of one model family over another may be explained by a different degree of training on the test task. To this end, we propose an effective method to adjust for the effect of training on the test task on benchmark evaluations. Put simply, to fine-tune each model under comparison on the same task-relevant data before evaluation. We then show that instances of emergent behavior disappear gradually as models train on the test task. Our work promotes a new perspective on the evaluation of large language models with broad implications for benchmarking and the study of emergent capabilities.222Code and fine-tuned models are available at https://github.com/socialfoundations/training-on-the-test-task
@@ 1 Introduction
In this work, we group strategies to utilize task knowledge at training time under the umbrella term of training on the test task. Examples of training on the test task include the use of instruction-tuning data or question answering templates during pre-training (Bai et al., 2023; StabilityAI, 2023; Groeneveld et al., 2024; Zhang et al., 2024). Models may also implicitly train on the test task when their pretraining data is selected through ablations on downstream benchmark evaluations (Gemma et al., 2024; MetaAI, 2024). We work from the premise that training on the test task is acceptable–or at least, unavoidable.
In a nutshell, we show that training on the test task strongly confounds model comparisons across different scales and model families. Moreover, it significantly obscures the study of emergent capabilities of large language models. Perhaps counterintuitively, we propose to mitigate the effects of training on the test task on benchmark evaluations by doing more of it. We show that we can effectively level the playing field by giving each model the same, sufficient task-specific fine-tuning before evaluation. This adjustment restores cleaner log-linear scaling and makes capabilities predictable based on much smaller model scales.
@@ 1.1 Our contributions
We introduce the term training on the test task to group a growing repertoire of practices that utilize knowledge about evaluation tasks at training time. We study its impact on present-day benchmark evaluations by critically examining the performance improvements of recent language models. Our analysis spans 56 different language models and two major active benchmarks, MMLU and GSM8K.
@@
We start in Section 2 by dividing models into those trained before November 2023 and those trained after. We find that for the same amount of pretraining compute, newer models strongly outperform older ones, on average by 7 percentage points in MMLU and 19 points in GSM8K. We then fine-tune all models on the same amount of task-specific data before evaluation. After fine-tuning on the same task data, newer models no longer outperform older ones. Rather, their performance equalizes. See Figure 1. This outcome suggests that the main difference between newer and older models is the extent to which they train on the test task.
@@
We propose a simple and effective method to adjust for the effect of training on the test task on benchmark evaluations. Put simply, to fine-tune each model on the same, sufficient amount of task-specific data before evaluation. To validate our method, we demonstrate its effectiveness in a controlled setting: we take the older models and fine-tune them on the test task. Remarkably this recreates the kind of performance differences observed between newer and older models, further suggesting that training on the test task explains the improvements of newer models. We then show that we can undo the advantage of the fine-tuned models over the other models by further fine-tuning all models on the test task (Section 3.1, Figure 3).
@@
Next, we give evidence that training on the test task may be a more dominant factor in benchmark performance than data contamination. To argue this point, we consider ARC and HellaSwag. Here, at first there appears to be no sign of newer models have an unfair advantage over older models. But after reformulating these benchmarks as MMLU-style multiple choice question answering tasks (MCQA), we see the same confounded results as for MMLU (Section 3.2, Figure 4). This suggest that the improvements of newer models on MMLU are likely not because of memorization of specific testing data, but rather due to an improved ability for MCQA tasks. Either way, our proposed adjustment recovers fair model comparisons.
@@
Then, we show how training on the test task distorts model family comparisons. Certain model families appear markedly superior to others before adjusting for test task training but not after adjustment (Section 4.1, Figure 6). We then demonstrate how training on the test task has inflated the perceived progress made by recent model families. After adjusting for its effect, newer models offer only modest improvements to the Pareto frontier of model performance relative to pre-training compute (Section 4.2, Figure 7).
@@
Finally, we demonstrate that training on the test task has profound implications for the study of emergent capabilities. The phenomenon of emergence disappears gradually as the amount of training on the test task grows (Section 5). Specifically, we can make capabilities visible and predictable from much smaller model scales, recovering cleaner log linear-scaling. Importantly, our adjustment also works in cases, like MMLU, where previous purported explanations of emergence, such as the choice of evaluation metric, do not suffice.
@@
Our work calls for a major reorientation of large language model evaluation. Model comparisons and claims of emergence are strongly confounded by the choice of training data relative to the test tasks. Rather than scrambling to detect and disallow various forms of training on the test task, we propose to “fight fire with fire”. When comparing models with different pre-training data, our recommendation is to give each model the same sufficient amount of fine-tuning on task-relevant data prior to evaluation.
@@
Although deliberate training on the test task can level the playing field, our work shows that generally significant fine-tuning is needed before the playing field is actually level. This requirement poses additional computational burden on the side of the evaluator. In addition, sufficient task-relevant training data might be expensive to source or generally unavailable for many tasks. While this is not an issue when model trainers also lack access to such data, training on proprietary task data could be difficult to correct for.
@@ 2 Adjusting for training on the test task
We choose MMLU (Hendrycks et al., 2020) and GSM8K (Cobbe et al., 2021) as a case study for investigating training on the test task in active benchmarks. MMLU tests for world knowledge, whereas GSM8K tests multi-step mathematical reasoning. These two benchmarks are arguably the most influential of the 2022-2024 period under study111For example, GPT 4 (Achiam et al., 2023), Claude 3 (Anthropic, 2024), Gemini (Gemini et al., 2023) and Llama 3 (MetaAI, 2024) all report and prominiently highlight MMLU and GSM8K performance. They are also included in the HuggingFace (HF) Open LLM Leaderboard v1 (Beeching et al., 2023), a popular leaderboard that evaluates and ranks models with publicly available weights. We evaluate models using LM Evaluation Harness (EleutherAI, 2024), in identical fashion to the HF leaderboard222Similarly to the HF leaderboard, we evaluate MMLU and GSM8K 5-shot, ARC 25-shot, and HellaSwag 10-shot..
We evaluate 56 base models, ranging in size from 70M to 70B parameters. See Appendix A.1 for the full list. The HF leaderboard’s FAQ makes the distinction between “base pretrained models” and instruction-tuned or chat models, arguing that this is necessary to ensure fair model comparisons. We select models that are categorized as “pretrained”. We check that the technical report of each of the selected models makes no mention of the model being fine-tuned. We only consider models for which the number of training tokens is known. This allows us to estimate the total amount of pretraining compute in FLOPs as , where is pretraining compute, is the number of model parameters, and is the number of training tokens.
While we focus primarily on MMLU and GSM8K due to their prominence, we find that the issue of training on the test task extends beyond these two benchmarks. Specifically, in Appendix D, we evaluate and discuss the impact of training on the test task for five additional benchmarks: MMLU Pro (Wang et al., 2024), GPQA (Rein et al., 2023), BBH (Suzgun et al., 2023), MuSR (Sprague et al., 2023), and MATH Level 5 (Hendrycks et al., 2021), which form the OpenLLM Leaderboard v2 (Fourrier et al., 2024a). Furthermore, in Appendix F we conduct similar experiments for an additional 36 instruction and chat models. We observe that our findings generalize remarkably well to instruction and chat models.
@@
We evaluate models on MMLU and GSM8K, and plot benchmark accuracy against pretraining compute in Figure 1 top. We observe that performance correlates with pretraining compute for both benchmarks. However, on the surface it appears that models released after November 2023 better leverage pretraining compute. In other words, for a given compute budget newer models are able to attain better benchmark performance.
We selected November 2023 as the temporal cutoff for our analysis because the technical reports of models released from late 2023 onward start referencing certain pre-training practices that may amount to training on test task. For example, Qwen (Bai et al., 2023), Olmo 1.7 (Groeneveld et al., 2024) and MAP Neo (Zhang et al., 2024) explicitly include instruction data during pretraining. StableLM 2 (StabilityAI, 2023) reformulates some of its pretraining datasets to better resemble downstream tasks such as question-answering. More subtly, the pretraining data mixtures of Gemma (Gemma et al., 2024) and Llama 3 (MetaAI, 2024) were determined through extensive ablations on downstream benchmark evaluations. We validate that our findings are robust to adjusting the temporal cutoff by a few months; see Appendix C.1 for details. Choosing specifically the month of November as the cutoff is therefore not critical for our analysis.
This raises an important question: Do newer models outperform older ones mainly because newer models trained more on the test task? At first sight, an answer seems elusive. After all, the pretraining data of most recent models is not publicly available. Retraining all models with the same training data and compute budget would be both infeasible and cost prohibitive. Nevertheless, in the next section, we propose a way to get at the answer by adjusting for the effect of training on the test task.
@@ 2.1 Adjusting for training on the test task by training on the test task
We propose to adjust for differences in test task training by fine-tuning all models on the same, sufficient amount of task-specific data before evaluation. To do so, we need a source of task-specific data for each of the tasks we consider. For multiple choice questioning answering, we use the auxiliary training set accompanying the HF MMLU repository333https://huggingface.co/datasets/cais/mmlu. This training set is not an i.i.d. split of MMLU. Instead, it consists of the training sets from other multiple-choice question-answering benchmarks, comprising approximately 100,000 training examples and 30 million tokens. For mathematical reasoning, we combine MetaMathQA (Yu et al., 2023b) and Orca-Math (Mitra et al., 2024), totalling approximately 600,000 training examples and 200M tokens. We fine-tune models for three epochs using standard hyperparameter choices, see Appendix A.2. The amount of compute required for fine-tuning is minimal compared to models’ pretraining compute.
We plot model scores on MMLU and GSM8K after fine-tuning in Figure 1 (bottom). We observe that after fine-tuning on task relevant data, both newer and older models follow remarkably similar scaling trends. That is, newer models no longer appear to outperform older models.
Remarkably, we observe that older models tend to benefit much more from fine-tuning on task-relevant data compared to newer models, see Figure 2. The improvements in older models are striking, often leaping from random chance accuracy to double-digit gains in accuracy. In contrast, fine-tuning provides comparatively little benefit to newer models. This observation suggests that newer models have already been exposed to a substantial amount of task-relevant data, making additional fine-tuning less impactful.
A potential concern is that our observations might result from our fine-tuning hyperparameters being systematically more favorable to older models. We verify that this is not the case by conducting a robustness check on the fine-tuning hyperparameters, see Appendix A.3.
@@ 2.2 Quantifying performance differences between newer and older models
We draw inspiration from scaling laws (Kaplan et al., 2020) in how we model benchmark accuracy to scale log-linearly with pretraining compute . To account for emergence (Wei et al., 2022b), we assume that models perform at the task’s random chance accuracy up to scaling to some point of emergence . We let the variable denote whether a model was trained after November 2023, and regress the model
where , and are the fit’s parameters, and is random noise. We focus on the coefficient , which corresponds to the average difference in benchmark performance between newer and older models after controlling for pretraining compute. We fit the model in Equation 1, and report the regression coefficient in Figure 1. We obtain R for all model fits.
Before adjusting for test task training, the estimated difference in performance between newer and older models are statistically significant, positive, and large. Specifically, recent models outperform older ones on average by over 7 accuracy points in MMLU and 19 accuracy points in GSM8K. These are remarkably large differences in benchmark performance. However, after the adjustment, the estimated coefficient is both small and not statistically significant. See Figure 1 bottom. That is, conditioned on all models training on the same amount of task-specific data, we find no evidence for a significant difference in benchmark performance between newer and older models.
Therefore, the performance of newer and older models equalizes when all models are exposed to the same amount of task-relevant data. This suggests that the impressive benchmark improvements of newer models are primarily attributable to newer models training more on the test task. We present a causal interpretation of results in Appendix B, outlying the assumptions necessary to establish a causal link between training on the test task and the benchmark improvements of newer models.
@@ 3 Recreating differences in benchmark performance
We have so far established that newer models strongly outperform older models for the same amount of pre-training compute. We now demonstrate how to recreate such differences in performance by actively manipulating how much models train on the test task. We do so in two ways. First, we fine-tune older models on task relevant data (Section 3.1). Second, we reformulate certain test tasks to use MMLU-style multiple choice prompts instead of “cloze” evaluations (Section 3.2). Both experiments recreate the kind of performance differences observed between newer and older models.
These results provide further evidence that the differences in performance between older and newer models are linked to test task training. They also demonstrate how test task training distorts benchmark evaluations. Fortunately, in both cases, we show that fine-tuning models on task-relevant data before evaluation is an effective mechanism for mitigating the bias introduced by training on the test task. In doing so, we systematically validate the proposed adjustment method.